1.    Given the following production function Q = 20L2K and the unit prices of labor and capital to be Birr 48 and 12 respectively, then

a. What combination of labor and capital maximizes output with a cost of Birr 720?

b. What is the maximum output?

c. Show the output maximizing condition graphically.




1
Expert's answer
2021-07-01T13:18:50-0400

a.

For solving this question, we will use Lagrange method

"Q = 20L^2K"

The equation for isocost line is "48L + 12K = 720"

So, using Lagrange method,

"L = 20L^2K - \lambda (48L + 12K - 720)"

Differentiating w.r.t L and K 

"20\times2LK - \lambda(48) = 0 ..................................... (1)"

"20L^2 - \lambda(12) = 0 ............................................. (2)"

Dividing equation (1) by (2),

"\frac{2K}{L }= \frac{48}{12}"

"K = 2L"

Putting this value of K in isocost line

"48L + 12(2L) = 720"

"72L = 720" "L = 10"

and"K = 2\times 10 = 20"

So, combination of labor that maximizes output is L = 10 and K = 20 units.


b.

Maximum output can be calculated by putting value of K and L in production function Q

So,

"Q = 20(10)^2(20) = 40000"

So, maximum output = 40000


c.

The graph below is showing the maximizing condition.




AB is the isocost line. Intercept on X axis is calculated by"\frac{720}{48} = 15" and on Y axis "\frac{720}{12} = 60" .

Q is the production function curve.

Point E is showing the point where we will have maximum production of 40000, using L = 10 and K = 20 units. 

The maximizing condition is that the slope of isocost line and production curve should be equal.